OpenStudy (anonymous):

|4x – 8| – 3x = 6x + 2 can anyone show me the steps to complete this equation ?

OpenStudy (anonymous):

At first it gets easier if you isolate the absolute expression, and separate it in two different inequalities

OpenStudy (anonymous):

oh, sorry, thats not an inequality

OpenStudy (anonymous):

|4x - 8| = 6x + 2 + 3x |4x - 8| = 9x + 2 4x - 8 = 9x + 2 or 4x - 8 = -(9x + 2) 4x - 9x = 2 + 8 --> 4x - 8 = -9x - 2 -5x = 10 --> 4x + 9x = -2 + 8 x = -2 -->13x = 6 x = 6/13